OpenStudy (anonymous):

Find an equation of the tangent line to the curve at the point (4π, 0). y = sin(sin(x))

OpenStudy (anonymous):

y = m x + b ( sin (sin (x) ) ' = cos (sin x) * cos (x) cos (sin 4 pi) * cos 4pi = cos ( 0 ) * 1 = 1*1 = 1 y = 1 * x + b y = x + b 0 = 4 pi + b b = - 4 pi y = x - 4pi